OpenStudy (anonymous):

A piece of string 25cm long goes exactly once round the edge of a plate in the form of a quadrant. Calculate the area of the quadrant. Take the value of pi to be 3.142.

OpenStudy (anonymous):

well i assume it is a quarter of a plate so what is the area of a quarter circle? whole circum.=2*pi*r so a quarter = (2*pi*r)/4=(pi*r)/2 =25(given) so pi*r=50....so r=50/pi put this in the area formula pi(50/pi)^2=2500/pi =796

OpenStudy (anonymous):